# The Sommerfeld-Rellich Framework for Scattering on Hyperbolic Space: Far-Field Patterns and Inverse Problems

### v2

@@ 6. Hyperbolic inverse obstacle scattering
where $\mathcal{B}$ denotes the boundary operator (Soft, hard or impedance obstacle). The existence and uniqueness of the scattering from obstacle has been proved in Theorem 4.6. Recall that we have shown that the total wave $u(x)$ can be written as
@@
(6.4) $\begin{cases}&-\Delta_{\mathbb{B}^{n}}u_{z,1}^{k}-\frac{(n-1)^{2}}{4}u_{z,1}^{k}=\mu^{2}u_{z,1}^{k},\ \ x\in\mathbb{B}^{n}\setminus\Omega_{1}\\ &u_{z,1}^{k}=0,\ \ x\in\partial\Omega_{1}\\ &\displaystyle{\frac{\partial(u_{z,1}^{k}-u^{i,k}_{z})}{\partial\rho}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh\left(\frac{\rho}{2}\right)(u_{z,1}^{k}-u^{i,k}_{z})=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right)},\end{cases}$
@@
(6.5) $\begin{cases}&-\Delta_{\mathbb{B}^{n}}u_{z,2}^{k}-\frac{(n-1)^{2}}{4}u_{z,2}^{k}=\mu^{2}u_{z,2}^{k},\ \ x\in\mathbb{B}^{n}\setminus\Omega_{2}\\ &u_{z,2}^{k}=0,\ \ x\in\partial\Omega_{2}\\ &\displaystyle{\frac{\partial(u_{z,2}^{k}-u^{i,k}_{z})}{\partial\rho}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh\left(\frac{\rho}{2}\right)(u_{z,2}-u^{i,k}_{z})=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right)}.\end{cases}$
@@
(6.6) $\begin{split}&u_{\infty,z,1}^{k,\mu}(\hat{x})\\ &\ \ =\left(\int_{0}^{\pi}(\sin t)^{-2\mu\mathrm{i}}dt\times\frac{A_{n,-\mu\mathrm{i}}}{2^{\frac{n-1}{2}}}\right)\left(\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot y+|y|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\frac{\partial u_{z,1}^{s,k}}{\partial\rho(y)}d\sigma_{\mathbb{B}^{n}}(y)\right.\\ &\ \ \ \ +\left.\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\partial}{\partial\rho(y)}\left(\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot\tanh(\frac{\rho(y)}{2})+|\tanh(\frac{\rho(y)}{2})|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\right)d\sigma_{\mathbb{B}^{n}}(y)\right)\end{split}$
@@
(6.7) $\begin{split}&u_{\infty,z,2}^{k,\mu}(\hat{x})\\ &\ \ =\left(\int_{0}^{\pi}(\sin t)^{-2\mu\mathrm{i}}dt\times\frac{A_{n,-\mu\mathrm{i}}}{2^{\frac{n-1}{2}}}\right)\left(\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot y+|y|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\frac{\partial u_{z,2}^{s,k}}{\partial\rho(y)}d\sigma_{\mathbb{B}^{n}}(y)\right.\\ &\ \ \ \ +\left.\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\partial}{\partial\rho(y)}\left(\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot\tanh(\frac{\rho(y)}{2})+|\tanh(\frac{\rho(y)}{2})|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\right)d\sigma_{\mathbb{B}^{n}}(y)\right)\end{split}$
@@ 7. Hyperbolic inverse potential scattering
with the far field pattern $u^{\mu}_{\infty}(\hat{x},\xi)$ given by
@@
The inverse medium problem in hyperbolic space addressed in this work concerns the simultaneous recovery of the medium $q(x)$ from the measurement of far-field pattern $u^{\mu}_{\infty}(\hat{x},\xi):$
@@
If we know the far-field pattern $u^{\mu}_{\infty}(\hat{x},\xi)$ for all incident direction $\xi$ and observable direction $\hat{x}$, then we can uniquely determine the potential $q(x)$.
@@
for all solutions $w_{1}$ and $w_{2}$ of (7.2) and (7.3). In order to prove that $q_{1}-q_{2}=0$, we need the following lemma (see [5]).

### v3

@@ 6. Hyperbolic inverse obstacle scattering
where $\mathcal{B}$ denotes the boundary operator (Soft, hard or impedance obstacle). The existence and uniqueness of the scattering from obstacle has been proved in Theorem 4.6. Recall that we have shown that the total wave field $u$ can be written as
@@
$\begin{cases}&-\Delta_{\mathbb{B}^{n}}u_{z,1}^{k}-\frac{(n-1)^{2}}{4}u_{z,1}^{k}=\mu^{2}u_{z,1}^{k},\ \ x\in\mathbb{B}^{n}\setminus\Omega_{1}\\ &u_{z,1}^{k}=0,\ \ x\in\partial\Omega_{1}\\ &\displaystyle{\frac{\partial(u_{z,1}^{k}-u^{i,k}_{z})}{\partial\rho}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh\left(\frac{\rho}{2}\right)(u_{z,1}^{k}-u^{i,k}_{z})=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right)},\end{cases}$
@@
$\begin{cases}&-\Delta_{\mathbb{B}^{n}}u_{z,2}^{k}-\frac{(n-1)^{2}}{4}u_{z,2}^{k}=\mu^{2}u_{z,2}^{k},\ \ x\in\mathbb{B}^{n}\setminus\Omega_{2}\\ &u_{z,2}^{k}=0,\ \ x\in\partial\Omega_{2}\\ &\displaystyle{\frac{\partial(u_{z,2}^{k}-u^{i,k}_{z})}{\partial\rho}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh\left(\frac{\rho}{2}\right)(u_{z,2}-u^{i,k}_{z})=o\left(\frac{1}{\sinh^{\frac{n-1}{2}}(\rho)}\right)}.\end{cases}$
@@
$\begin{split}&u_{\infty,z,1}^{k,\mu}(\hat{x})\\ &\ \ =\left(\int_{0}^{\pi}(\sin t)^{-2\mu\mathrm{i}}dt\times\frac{A_{n,-\mu\mathrm{i}}}{2^{\frac{n-1}{2}}}\right)\left(\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot y+|y|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\frac{\partial u_{z,1}^{s,k}}{\partial\rho(y)}d\sigma_{\mathbb{B}^{n}}(y)\right.\\ &\ \ \ \ +\left.\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\partial}{\partial\rho(y)}\left(\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot\tanh(\frac{\rho(y)}{2})+|\tanh(\frac{\rho(y)}{2})|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\right)d\sigma_{\mathbb{B}^{n}}(y)\right)\end{split}$
@@
$\begin{split}&u_{\infty,z,2}^{k,\mu}(\hat{x})\\ &\ \ =\left(\int_{0}^{\pi}(\sin t)^{-2\mu\mathrm{i}}dt\times\frac{A_{n,-\mu\mathrm{i}}}{2^{\frac{n-1}{2}}}\right)\left(\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot y+|y|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\frac{\partial u_{z,2}^{s,k}}{\partial\rho(y)}d\sigma_{\mathbb{B}^{n}}(y)\right.\\ &\ \ \ \ +\left.\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\partial}{\partial\rho(y)}\left(\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot\tanh(\frac{\rho(y)}{2})+|\tanh(\frac{\rho(y)}{2})|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\right)d\sigma_{\mathbb{B}^{n}}(y)\right)\end{split}$
@@ 7. Hyperbolic inverse potential scattering
with the far field pattern $u^{s}_{\infty}(\hat{x},\mu,\xi)$ given by
@@
The inverse medium problem in hyperbolic space addressed in this work concerns the simultaneous recovery of the medium $q(x)$ from the measurement of far-field pattern $u^{s}_{\infty}(\hat{x},\mu,\xi):$
@@
If we know the far-field pattern $u^{\mu}_{\infty}(\hat{x},\xi)$ for all incident direction $\xi$ and observable direction $\hat{x}$, then we can uniquely determine $q(x)$.
@@
for all solutions $w_{1}$ and $w_{2}$ of (7.2) and (7.3). In order to prove that $q_{1}-q_{2}=0$, we need the following lemma.